OpenStudy (kingj):

A line contains points (2, -2) and (-4, 1). Another line graphed in the same coordinate plane contains points (1, 2) and (-3, 4). Based on the slopes of these lines, are they parallel, perpendicular or neither?

OpenStudy (campbell_st):

slope of line 1 (2,-2) and (-4, 1) m = (1 - (-2))/((-4) - 2) m = 3/(-6) m = -1/2 2nd line (1,2)(-3, 4) m = (4 -2)/((-3) -1) m = 2/(-4) m - -1/2 the lines are parallel as they have the same gradient.